OpenStudy (anonymous):

From the Cartesian coordinates (2,-10), I want (r, theta). So far, I have that r=sqrt(104). For the angle, I took the fact sqrt(104)*sin(theta)=2/sqrt(104). Now, sin(theta)=2/104, giving theta=asin(1/52). That *seems* right. "Ding! Survey *SAYS!* Nope." :/

OpenStudy (jdoe0001):

well. your "r" is correct since \(\bf r=\sqrt{x^2+y^2}\) now to find the angle ( 2 , -10 ) x y thus \(\bf tan(\theta)=\cfrac{y}{x}\to \cfrac{-\cancel{ 10 }}{\cancel{ 2 }}\to -5\quad thus \\ \quad \\ tan^{-1}[tan(\theta)]=tan^{-1}(-5)\implies \theta=tan^{-1}(-5)\)

OpenStudy (anonymous):

Funny, that is *exactly* what I tried the first time. I should have mentioned, it's looking for answers with r>=0 and 0<theta<2pi.

OpenStudy (anonymous):

I got the same answer, fwiw. :)

OpenStudy (jdoe0001):

well... the "r" you found Is >=0 and the angle that gives will be within that range

OpenStudy (anonymous):

I thought so, too.

OpenStudy (jdoe0001):

I'm thinking is expecting a constant value

OpenStudy (anonymous):

I'm not sure I follow.

OpenStudy (jdoe0001):

is not expecting sqrt(...) operation or atan() operation is expecting a solid number for both

OpenStudy (anonymous):

Actually, it accepted atan(3/19) for the previous problem. Anyhow, I'll check in a little later...I've been neglecting the wife and kids for a big chunk of the day now. The little one is leaning on my leg, begging to go outside and play. Hard to say no to her. :)

OpenStudy (jdoe0001):

hehe, indeed

OpenStudy (anonymous):

I'm closing this one for now. Final is next week, and I'd rather not get stuck. There were already two problems on this assignment where I technically had the right answer, but the system wouldn't give me credit. Kind of annoying. :/ Thanks for your help!